# Indicator Functions Detect Tangentially Transient Behaviour on Decaying Normally Hyperbolic Invariant Manifolds

### v1

@@ Abstract
We study the decay scenario of a codimension-2 NHIM in a three degrees of freedom Hamiltonian system under increasing perturbation when the NHIM loses its normal hyperbolicity. On one hand, we follow this decay in the Poincaré map for the internal dynamics of the NHIM. On the other hand, we also follow the decay in a time delay function calculated on a 2-dimensional plane in the phase space of the system. In addition, we observe the role of tangential transient effects on the decaying NHIM and their manifestation in the delay time indicator function. Thereby we obtain ideas on how the decay of NHIMs and the tangential transient effects are encoded in indicator functions. As an example of demonstration, we use the motion of an electron in a perturbed magnetic dipole field.
@@ 1 Introduction
The global dynamics of a chaotic dynamical system is governed to a large extent by its unstable invariant subsets in the phase space, in particular by the ones of codimension 2. These subsets have stable and unstable manifolds of codimension 1 which can form walls and tubes in the phase space which direct the general dynamics. These considerations of codimensions hold equally well for the constant energy manifold of Hamiltonian flows as for the corresponding Poincaré maps. A remarkable kind of invariant sets in the phase space are normally unstable invariant surfaces known under the name “Normally Hyperbolic Invariant Manifolds”, standard abbreviation NHIMs. For general properties of these manifolds see [1].
@@
For any important subsets in a phase space, the question of their stability always arrises under general perturbations of the system. For NHIMs, the following persistence theorem holds: NHIMs survive perturbations and keep their global topology as long as their normal instability remains larger than their tangential instability. This property is called “normal hyperbolicity”. Several methods to prove the persistence theorem of NHIMs are developed in the references [1, 2, 3, 4]. Examples indicate that NHIMs start to decay locally as soon as in some point on the NHIM the tangential instability becomes larger than the normal instability. For three instructive examples see Figs. 4 and 5 and their explanation in [2], example 1.1 with Fig.1.3 in subsection 1.2.1 of [4], and Fig.9 and its explanation in section 6 of [5].
@@
The persistence theorem does not imply the structural stability of the internal dynamics of the NHIMs. Under general perturbations, we find qualitative changes of the internal dynamics and in particular bifurcations of the important internal periodic orbits within the NHIMs. These changes also include the creation of chaos in the internal dynamics. However, according to the persistence theorem the NHIM survives and keeps its topology as long as the normal hyperbolicity property is conserved. The development scenario of the internal dynamics of a NHIM under any perturbation can be investigated and presented graphically best by the restriction of the Poincaré map to the NHIM.
@@
In the following, we are interested in Hamiltonian systems with 3 degrees of freedom (3-dof). They have a complete phase space of dimension 6, a constant energy manifold of dimension 5, and Poincaré maps acting on domains of dimension 4. Therefore, the most interesting invariant subsets in the constant energy manifold are the ones of dimension 3, and they correspond to invariant surfaces of dimension 2 in the domain of the Poincaré map. In most cases, it is simpler to argue for maps than for flows. Therefore in this article, we will present our ideas mainly for maps. For 3-dof systems the codimension-2 NHIMs in the map are 2-dimensional and then the map of their internal dynamics can be displayed by 2-dimensional graphics. We use this graphical representation extensively in the present article.
@@
Some investigations of the loss of normal hyperbolicity of NHIMs have been reported in [6, 7, 8, 9]. These examples either use 2-dof systems or use simplifications which are equivalent to a reduction of the problem to 2 degrees of freedom. They are mainly concerned with the implications of the loss of normal hyperbolicity to transition state theory and transport problems. They are less concerned with the remnants of the NHIMs during their decay to fragments which are partially of lower dimensions, a problem with which we are concerned in the present article. For further interesting publications on changes of NHIMs and their invariant manifolds under parameter changes see also [10, 11, 12, 13, 14, 15, 16].
@@
In the present article, we deal with the problem of the disintegration of NHIMs when they lose their normal hyperbolicity and with the problem of how we observe this decay. We investigate the remaining fragments numerically with the help of phase space structure indicator functions. Therefore let us give here some basic introductory remarks on indicator functions.
@@
When a 0-dimensional unstable invariant set, i.e. a hyperbolic fixed point of a map, collides with a stable fixed point and runs through a saddle-center bifurcation then it disappears completely from the domain of the map. Using the normal form approximation of the map around the hyperbolic equilibrium point, we can observe how these fixed points disappear into the complex generalization of the domain of the map.
@@
First, we consider a set of parameter values where the unstable fixed point still exists and let us discuss what we observe in indicator functions. The most appropriate indicator function for this purpose is the time delay function which indicates the time during which a general trajectory stays in some neighbourhood of the unstable fixed point. If the trajectory starts exactly on the stable manifold of the fixed point, then this trajectory converges to the fixed point, it stays in the neighbourhood of the fixed point for an infinite future and accordingly, the time delay diverges. When it starts not exactly on the stable manifold but close to it, then this trajectory comes close to the fixed point, stays in its neighbourhood for some time but leaves this neighbourhood again after a finite time and accordingly, the time delay is large but finite.
@@
If we plot this time delay over some surface of initial conditions in the phase space then we detect the intersections between the stable manifold of the fixed point and this surface. This basic idea also works for the detection of stable manifolds of higher dimensional unstable invariant subsets and it is the basic idea of how we search for NHIMs numerically, for all the details see [17].
@@
For a 0-dimensional unstable fixed point, the invariant set disappears at once completely in a single bifurcation at one specific value of the perturbation parameter. For higher dimensional invariant subsets the decay of the invariant set is a process which starts at a particular value of the perturbation parameter. Then the decay continues over some possibly large interval of the perturbation parameter. Parts of the original 2-dimensional NHIM surface remain while other parts decay and leave behind only a ( in general fractal ) collection of fragments which partially have lower dimensions than the original NHIM. In the most extreme case, the NHIM in the map turns into a fractal powder only, for an example of such a scenario see [5].
@@
In contrast look at decaying parts. Here the NHIM surface decays into lower than 2-dimensional parts, let us look at a particular part with a ( maybe fractal ) dimension $k$. Their stable manifolds have dimension $k+1$. The transverse intersection of these stable manifolds with the 2-dimensional intersection surface over which we plot the indicator function has dimension $k-1$. In particular, if $k<1$ then this intersection is empty in general. That is, we do not see any corresponding singularities in the plot.
However, when we start in regions close to some remnants of the stable manifolds of the NHIM then there is some large but finite time delay compared with other regions nearby. The effect looks as if the infinite singularities of the indicator function are removed and only a high background is left. If there is a fractal collection of remnants, then the time delay function has a lot of finite maxima which gives this function a complicated appearance. For a previous observation of this type of complicated behaviour in a time delay function along a 1-dimensional line of initial conditions see [18].
@@
Finally, we consider a general trajectory which approaches the neighbourhood of the NHIM by motion close to its surviving parts of the stable manifold. Then this trajectory in the neighbourhood of the decaying NHIM has two competing possibilities to leave the neighbourhood of the surviving parts of the NHIM again. First, it can leave along the unstable manifolds of the surviving parts of the NHIM. Second, it can move mainly tangentially to the NHIM over the boundary of decay and leave the surviving parts of the NHIM along this direction. Or more generally it can leave by a combination of these two routes.
In this article, we treat the case of 3-dof systems and NHIMs of dimension 2 in the map. When such a NHIM starts to decay then we find surviving parts of 2-dimensional surfaces. In the general case, the internal dynamics of such parts contain KAM curves and chaos layers. Then the interior of KAM curves ( and this includes chaos layers inside ) are trapped on these surviving parts of the NHIM forever.
As a short side remark let us mention briefly the essential differences which we find for systems with even more degrees of freedom, let us say $k>3$ in number. Then the dimension of the domain of the map is $22k-$ 2. Now assume that we have a codimension 2 NHIM, i.e. it has dimension $22k-4>$ 2 in the map. There are still KAM curves of the internal dynamics of the NHIM. And we have chaos layers. However, for $k>2$, we have Arnold diffusion in the internal dynamics of the NHIM, all the chaos layers are interconnected and a general trajectory starting in some fine chaos layer in the long run can diffuse over the whole chaotic part of the internal dynamics.
@@
In this work, we take the motion of a charged particle in a perturbed magnetic dipole field as an example of demonstration. The perturbation is the addition of a quadrupole contribution to the magnetic field. This example has one unusual property. Namely, the decay of the NHIM starts as soon as the perturbation parameter, and the magnitude of the quadrupole contribution becomes different from 0. So we do not have some interval of the perturbation parameter, where the NHIM persists in the topology which it had for perturbation 0. However, this is no problem for the topic of the present article, since we are interested in the decay of NHIMs anyway. Otherwise, this magnetic dipole example is an excellent and typical example to study the decay scenario of a NHIM. Details of our previous studies of this magnetic dipole example and many further references for this system can be found in [20].
The present article is organised as follows: In section 2, we present the basic dynamics of our example of demonstration, in particular, the construction of the main NHIM and its stable and unstable manifolds. In section 3, we show the plots of the indicator function, i.e. the time delay function, and for comparison also the corresponding plots of the Poincaré map restricted to the NHIM. Section 4 contains our conclusions and some final remarks.
@@ 2 Example: The motion of an electrically charged particle around a perturbed magnetic dipole
Here $A_{0}$ is the magnitude of the magnetic dipole moment which is aligned in the $z$ direction. The term in Eq.1 containing $A_{0}$ is the magnetic dipole potential. And $\epsilon$ is the magnitude of the quadrupole perturbation of the field, which serves as a perturbation parameter in the investigation of the development scenario of the NHIM $M$ of this particular system. The term in Eq.1 containing $\epsilon$ is the magnetic potential of one particular quadrupole component. The phase space of the full system is 6-dimensional and has the coordinates $(r,p_{r},\phi,p_{\phi},z,p_{z})$.
The dipole potential has a rotational symmetry around the $z$-axis and correspondingly under the influence of the dipole field only, i.e. in the unperturbed case $\epsilon=0$ the $z$-component of the angular momentum of the particle, i.e. $L_{z}=p_{\phi}$, is conserved and it acts as if it would be a parameter of the system. In this case, the system reduces to a 2-dof system. For the quadrupole field, we have chosen a component which is not symmetric around the $z$-axis such that for $\epsilon\neq 0$ the angular momentum $L_{z}$ is no longer conserved. In this case, the system can no longer be reduced to a 2-dof system.
In the next section, we show numerical results for the perturbation of the system where we start with the unperturbed case of $\epsilon=0$ and then increase $\epsilon$ slowly to observe the effects of the perturbation away from the partially integrable case. Let us denote as $M_{0}$ the NHIM of the system for the unperturbed case and $M_{\epsilon}$ the NHIM for the perturbed case. In particular, we will study the change of the NHIM $M_{\epsilon}$ of the system under this perturbation. Therefore, it is important to understand the structure of the NHIM $M_{0}$ well in the unperturbed case. For $\epsilon=0$, the 6-dimensional phase space of the complete 3-dof system foliates into a continuum of 2-dof systems, one for each possible value of the conserved quantity $L_{z}$. The phase space of each one of these reduced 2-dof systems is a copy of the 4-dimensional space of the variables $(r,p_{r},z,p_{z})$.
@@
Having this pile construction in mind it is easy to imagine the structure of the NHIM $M_{0}$ of the unperturbed full 3-dof system. First, let us consider one particular value of the total energy $E$. This cuts out a 5-dimensional surface of the full phase space and a corresponding pile of 3-dimensional energy surfaces of the reduced 2-dof systems.
@@
This pile construction also induces a pile construction of the Poincaré map and a foliation of its 4-dimensional domain into 2-dimensional leaves belonging to the various values of $L_{z}$. Now imagine that the reduced Poincaré map has a hyperbolic fixed point which exists for some interval of $L_{z}$ values. In the pile construction this continuum of fixed points forms a continuous line of points and by the formation of the Cartesian product with the circle representing the cyclic angle $\phi$ it forms a 2-dimensional surface with the topology of a cylinder segment. This surface is invariant by construction and it inherits normal hyperbolicity from the hyperbolicity of the fixed point in the reduced system. Thereby this surface is the NHIM $M^{P}_{0}$ of the unperturbed partially integrable system in the 4-dimensional domain of its Poincaré map.

### v2

@@ Abstract
We study the decay scenario of a codimension-2 NHIM in a three-degrees-of-freedom Hamiltonian system under increasing perturbation when the NHIM loses its normal hyperbolicity. On one hand, we follow this decay in the Poincaré map for the internal dynamics of the NHIM. On the other hand, we also follow the decay in a time delay function calculated on a 2-dimensional plane in the phase space of the system. In addition, we observe the role of tangential transient effects on the decaying NHIM and their manifestation in the delay time indicator function. We obtain ideas on how the decay of NHIMs and the tangential transient effects are encoded in indicator functions. As an example of demonstration, we use the motion of an electron in a perturbed magnetic dipole field.
@@ 1 Introduction
The global dynamics of a chaotic dynamical system is governed to a large extent by its unstable invariant subsets in the phase space, in particular by the ones of codimension 2. These subsets have stable and unstable manifolds of codimension 1 which can form walls and tubes in the phase space that direct the general dynamics. These considerations of codimensions hold equally well for the constant energy manifold of Hamiltonian flows as well as for the corresponding Poincaré maps. A remarkable kind of invariant sets in the phase space are normally unstable invariant surfaces known under the name “Normally Hyperbolic Invariant Manifolds”, standard abbreviation NHIMs. For general properties of these manifolds see [1].
@@
For any important subsets in a phase space, the question of their stability always arises under general perturbations of the system. For NHIMs, the following persistence theorem holds: NHIMs survive perturbations and keep their global topology as long as their normal instability remains larger than their tangential instability. This property is called “normal hyperbolicity”. Several methods to prove the persistence theorem of NHIMs are developed in the references [1, 2, 3, 4]. Examples indicate that NHIMs start to decay locally as soon as in some point on the NHIM the tangential instability becomes larger than the normal instability. For three instructive examples see Figs. 4 and 5 and their explanation in [2], example 1.1 with Fig.1.3 in subsection 1.2.1 of [4], and Fig.9 and its explanation in section 6 of [5].
@@
The persistence theorem does not imply the structural stability of the internal dynamics of the NHIMs. Under general perturbations, we find qualitative changes in the internal dynamics and in particular bifurcations of the important internal periodic orbits within the NHIMs. These changes also include the creation of chaos in the internal dynamics. However, according to the persistence theorem, the NHIM survives and keeps its topology as long as the normal hyperbolicity property is conserved. The development scenario of the internal dynamics of a NHIM under any perturbation can be investigated and presented graphically best by the restriction of the Poincaré map to the NHIM.
@@
In the following, we are interested in Hamiltonian systems with 3 degrees of freedom (3-dof). They have a complete phase space of dimension 6, a constant energy manifold of dimension 5, and Poincaré maps acting on domains of dimension 4. Therefore, the most interesting invariant subsets in the constant energy manifold are the ones of dimension 3, and they correspond to invariant surfaces of dimension 2 in the domain of the Poincaré map. In most cases, it is simpler to argue for maps than for flows. Therefore in this article, we will present our ideas mainly for maps. For 3-dof systems, the codimension-2 NHIMs in the map are 2-dimensional and then the map of their internal dynamics can be displayed by 2-dimensional graphics. We use this graphical representation extensively in the present article.
@@
Some investigations of the loss of normal hyperbolicity of NHIMs have been reported in [6, 7, 8, 9]. These examples either use 2-dof systems or use simplifications which are equivalent to a reduction of the problem to 2 degrees of freedom. They are mainly concerned with the implications of the loss of normal hyperbolicity to transition state theory and transport problems. They are less concerned with the remnants of the NHIMs during their decay to fragments which are partially of lower dimensions, a problem with which we are concerned in the present article. For further interesting publications on changes of NHIMs and their invariant manifolds under parameter changes see also [10, 11, 12, 13, 14, 15, 16, 17].
@@
In the present article, we deal with the problem of the disintegration of NHIMs when they lose their normal hyperbolicity and with the problem of how we observe this decay. We investigate the remaining fragments numerically with the help of phase space structure indicator functions. Then let us give here some basic introductory remarks on indicator functions.
@@
When a 0-dimensional unstable invariant set, i.e. a hyperbolic fixed point of a map, collides with a stable fixed point and runs through a saddle-center bifurcation, then it disappears completely from the domain of the map. Using the normal form approximation of the map around the hyperbolic equilibrium point, we can observe how these fixed points disappear into the complex generalization of the domain of the map.
@@
First, we consider a set of parameter values where the unstable fixed point still exists. Let us discuss what we observe in indicator functions. The most appropriate indicator function for this purpose is the time delay function which indicates the time during which a general trajectory stays in some neighbourhood of a particular region, for example, an unstable fixed point. If the trajectory starts exactly on the stable manifold of the fixed point, then this trajectory converges to the fixed point, it stays in the neighbourhood of the fixed point for an infinite future and accordingly, the time delay diverges. When it starts not exactly on the stable manifold but close to it, then this trajectory comes close to the fixed point, stays in its neighbourhood for some time but leaves this neighbourhood again after a finite time and accordingly, the time delay is large but finite.
@@
If we plot this time delay over some surface of initial conditions in the phase space, then we detect the intersections between the stable manifold of the fixed point and this surface. This basic idea also works for the detection of stable manifolds of higher-dimensional unstable invariant subsets, and it is the basic idea of how we search for NHIMs numerically. For all the details, see [18].
@@
For a 0-dimensional unstable fixed point, the invariant set disappears at once completely in a single bifurcation at one specific value of the perturbation parameter. For higher dimensional invariant subsets, the decay of the invariant set is a process which starts at a particular value of the perturbation parameter. Then the decay continues over some possibly large interval of the perturbation parameter. Parts of the original 2-dimensional NHIM surface remain while other parts decay and leave behind only a ( in general fractal ) collection of fragments which partially have lower dimensions than the original NHIM. In the most extreme case, the NHIM in the map turns into a fractal powder only; for an example of such a scenario see [5].
@@
In contrast, look at decaying parts. Here the NHIM surface decays into lower than 2-dimensional parts. Let us look at a particular part with a ( maybe fractal ) dimension $k$. Their stable manifolds have dimension $k+1$. The transverse intersection of these stable manifolds with the 2-dimensional intersection surface over which we plot the indicator function has dimension $k-1$. In particular, if $k<1$ then this intersection is empty in general. That is, we do not see any corresponding singularities in the plot.
However, when we start in regions close to some remnants of the stable manifolds of the NHIM then there is some large but finite time delay compared with other regions nearby. The effect looks as if the infinite singularities of the indicator function are removed and only a high background is left. If there is a fractal collection of remnants, then the time delay function has a lot of finite maxima which gives this function a complicated appearance. For a previous observation of this type of complicated behaviour in a time delay function along a 1-dimensional line of initial conditions, see [19].
@@
Finally, we consider a general trajectory that approaches the neighbourhood of the NHIM by motion close to its surviving parts of the stable manifold. Then this trajectory in the neighbourhood of the decaying NHIM has two competing possibilities to leave the neighbourhood of the surviving parts of the NHIM again. First, it can leave along the unstable manifolds of the surviving parts of the NHIM. Second, it can move mainly tangentially to the NHIM over the boundary of decay and leave the surviving parts of the NHIM along this direction. Or more generally, it can leave by a combination of these two routes.
In this article, we treat the case of 3-dof systems and NHIMs of dimension 2 in the map. When such a NHIM starts to decay, then we find surviving parts of 2-dimensional surfaces. In the general case, the internal dynamics of such parts contain KAM curves and chaos layers. Then the interior of KAM curves ( and this includes chaos layers inside ) are trapped on these surviving parts of the NHIM forever.
As a short side remark let us mention briefly the essential differences which we find for systems with even more degrees of freedom, let us say $k>3$ in number. Then the dimension of the domain of the map is $22k-$ 2. Now assume that we have a codimension 2 NHIM, i.e. it has dimension $22k-4>$ 2 in the map. There are still KAM curves of the internal dynamics of the NHIM. And we have chaos layers. However, for $k>3$, we have Arnold diffusion in the internal dynamics of the NHIM, all the chaos layers are interconnected and a general trajectory starting in some fine chaos layer in the long run can diffuse over the whole chaotic part of the internal dynamics.
@@
In this work, we take the motion of a charged particle in a perturbed magnetic dipole field as an example of demonstration. The perturbation is the addition of a quadrupole contribution to the magnetic field. This example has one unusual property. Namely, the decay of the NHIM starts as soon as the perturbation parameter, i.e. the magnitude of the quadrupole contribution becomes different from 0. So we do not have some interval of the perturbation parameter, where the NHIM persists in the topology which it had for perturbation 0. However, this is no problem for the topic of the present article, since we are interested in the decay of NHIMs anyway. Otherwise, this magnetic dipole example is an excellent and typical example to study the decay scenario of a NHIM. Details of our previous studies of this magnetic dipole example and many further references for this system can be found in [21].
The present article is organised as follows: In section 2, we present the basic dynamics of our example of demonstration, in particular, the construction of the main NHIM and its stable and unstable manifolds. In section 3, we give a short explanation of our method to search for and represent the Poincaré map of the internal dynamics of the NHIMs. In section 4, we show the plots of the indicator function, i.e. the time delay function, and for comparison also the corresponding plots of the Poincaré map restricted to the NHIM. The last section contains our conclusions and final remarks.
@@ 2 Example: The motion of an electrically charged particle around a perturbed magnetic dipole
Here $A_{0}$ is the magnitude of the magnetic dipole moment which is aligned in the $z$-direction. The term in Eq.1 containing $A_{0}$ is the magnetic dipole potential. And $\epsilon$ is the magnitude of the quadrupole perturbation of the field, which serves as a perturbation parameter in the investigation of the development scenario of the NHIM of this particular system. The term in Eq.1 containing $\epsilon$ is the magnetic potential of one particular quadrupole component. The phase space of the full system is 6-dimensional and has the coordinates $(r,p_{r},\phi,p_{\phi},z,p_{z})$. Let us consider a fixed value of the total energy $E$ for the next considerations and denote as $M_{0}$ the NHIM of the system for the unperturbed case and $M_{\epsilon}$ the NHIM for the perturbed case.
The dipole potential has a rotational symmetry around the $z$-axis and correspondingly under the influence of the dipole field only, i.e. in the unperturbed case $\epsilon=0$ the $z$-component of the angular momentum of the particle, i.e. $L_{z}=p_{\phi}$, is conserved and it acts as a parameter of the system. In this case, the total 3-dof system reduces to a 2-dof system with an effective potential energy is given by
$V_{eff}(\rho,z)=\frac{1}{2m}\left(\frac{L_{z}}{r}-A_{0}\frac{r}{(r^{2}+z^{2})^{3/2}}\right)^{2}.$ (2)
This 2-dimensional effective potential energy has a saddle point which generates a Lyapunov family of unstable hyperbolic periodic orbits $\gamma_{L_{z}}$ in the phase space of the 2-dof system when we change the value of $L_{z}$ [18, 21]. Those hyperbolic periodic orbits parametrized by the $L_{z}$ are the basic constructing blocks of the NHIM in the phase space of the 3-dof symmetric unperturbed system. To construct the NHIM $M_{0}$ for this system, let us consider the union of all the hyperbolic periodic orbits in the Lyapunov family and take the Cartesian product with the circle $S^{1}$ that represents the angle $\phi$. Then the NHIM $M_{0}$ of the 3-dof unperturbed symmetric system is
$M_{0}=\bigcup_{L_{z}}\gamma_{L_{z}}\times S^{1}.$ (3)
In the same way, we construct the stable and unstable manifolds $W^{s/u}(M_{0})$ of the NHIM $M_{0}$ taking the union of all the stable and unstable manifolds $W^{s/u}(\gamma_{L_{z}})$ of the hyperbolic periodic orbits $\gamma_{L_{z}}$ and the Cartesian product with the circle $S^{1}$,
$W^{s/u}(M_{0})=\bigcup_{L_{z}}W^{s/u}(\gamma_{L_{z}})\times S^{1}.$ (4)
For the quadrupole field, we have chosen a component which is not symmetric around the $z$-axis such that for $\epsilon\neq 0$ the angular momentum $L_{z}$ is no longer conserved. In this case, the system can no longer be reduced to a 2-dof system.
In the next section, we show numerical results for the perturbation of the system where we start with the unperturbed case of $\epsilon=0$ and then increase $\epsilon$ slowly to observe the effects of the perturbation away from the partially integrable case.
In particular, we will study the change of the NHIM $M_{\epsilon}$ of the system under this perturbation. Therefore, it is important to understand the structure of the NHIM $M_{0}$ well in the unperturbed case. For $\epsilon=0$, the 6-dimensional phase space of the complete 3-dof system foliates into a continuum of 2-dof systems, one for each possible value of the conserved quantity $L_{z}$. The phase space of each one of these reduced 2-dof systems is a copy of the 4-dimensional space of the variables $(r,p_{r},z,p_{z})$.
@@
Having this pile construction in mind it is easy to imagine the structure of the NHIM $M_{0}$ of the unperturbed full 3-dof system. First, let us consider one particular value of the total energy $E$. This cuts out a 5-dimensional manifold of the full phase space and a corresponding pile of 3-dimensional energy manifold of the reduced 2-dof systems.
@@
This pile construction also induces a pile construction of the Poincaré map and a foliation of its 4-dimensional domain into 2-dimensional leaves belonging to the various values of $L_{z}$. Now imagine that the reduced Poincaré map has a hyperbolic fixed point $X_{L_{z}}$ which exists for some interval of $L_{z}$ values. In the pile construction, this continuum of fixed points forms a continuous line of points and by the formation of the Cartesian product with the circle $S^{1}$ representing the cyclic angle $\phi$ it forms a 2-dimensional manifold with the topology of a cylinder segment. We denote the NHIM in the Poincaré map as $M^{P}_{0}$. Then, the NHIM $M^{P}_{0}$ is given by
$M^{P}_{0}=\bigcup_{L_{z}}X_{L_{z}}\times S^{1}.$ (5)
This manifold is invariant under the Poincaré map by construction, and it inherits normal hyperbolicity from the hyperbolicity of the fixed point in the reduced system. Thereby this surface is the NHIM $M^{P}_{0}$ of the unperturbed partially integrable system in the 4-dimensional domain of its Poincaré map. In an analogous way, we can construct the stable and unstable manifolds of the NHIM $M^{P}_{0}$ corresponding to the Poincaré map as
$W^{s/u}(M^{P}_{0})=\bigcup_{L_{z}}W^{s/u}(X_{L_{z}})\times S^{1}.$ (6)